OpenStudy (anonymous):

Solve the equation in the interval [0, 2pi]. (Cos (x))^2=1/9

OpenStudy (anonymous):

Solve the equation in the interval [0, 2pi]. (Cos (x))^2=1/9

OpenStudy (anonymous):

Solve the equation in the interval [0, 2pi]. (Cos (x))^2=1/9